Exotic Food Inc., Capital Budgeting Case CASE SUMMARY Exotic Food Inc., a food processing company located in Herndon, VA, is considering adding a new division to produce fresh ginger juice. Following the ongoing TV buzz about significant health benefits derived from ginger consumption, the managers believe this drink will be a hit. However, the CEO questions the profitability of the venture given the high costs involved. To address his concerns, you have been asked to evaluate the project using three capital budgeting techniques (i.e., NPV, IRR and Payback) and present your findings in a report. CASE OVERVIEW The main equipment required is a commercial food processor which costs $200,000. The shipping and installation cost of the processor from China is $50,000. The processor will be depreciated under the MACRS system using the applicable depreciation rates are 33%, 45%, 15%, and 7% respectively. Production is estimated to last for three years, and the company will exit the market before intense competition sets in and erodes profits. The market value of the processor is expected to be $100,000 after three years. Net working capital of $2,000 is required at the start, which will be recovered at the end of the project. The juice will be packaged in 20 oz. containers that sell for $3.00 each. The company expects to sell 150,000 units per year; cost of goods sold is expected to total 70% of dollar sales. Weighted Average Cost of Capital (WACC): Exotic Food’s common stock is currently listed at $75 per share; new preferred stock sells for $80 per share and pays a dividend of $5.00. Last year, the company paid dividends of $2.00 per share for common stock, which is expected to grow at a constant rate of 10%. The local bank is willing to finance the project at 10.5% annual interest. The company’s marginal tax rate is 35%, and the optimum target capital structure is: Common equity 50% Preferred 20% Debt 30% Your main task is to compute and evaluate the cash flows using capital budgeting techniques, analyze the results, and present your recommendations whether the company should take on the project. What is the Weighted Average Cost of Capital (WACC)? Compute the after-tax cost of debt Compute the cost of common equity Compute the cost of preferred stock Compute the Weighted Average Cost of Capital (WACC) Using a WACC of 15%, apply four capital budgeting techniques to evaluate the project, assuming the Free Cash Flows are as follows: Years Free Cash Flows 0 $ -252,000.00 1 $118,625.00 2 $127,125.00 3 $181,000.00 The four techniques are NPV, IRR, MIRR, and discounted Payback. Assume the reinvestment rate to be 8% for the MIRR. Also, assume that the business will only accept projects with a payback period of two and half years or less.
Answer #1

What is the Weighted Average Cost of Capital (WACC)?

Compute the after-tax cost of debt

Pre tax cost of debt = 10.5%

Tax rate, T = 35%

Post tax cost of debt, Kd = Pre tax cost of debt x (1 - T) = 10.5% x (1 - 35%) = 6.825%

---------------------------

Compute the cost of common equity

Last year's dividend, D0 = 2

Dividend growth rate, g = 10%

Expected dividend next year, D1 = D0 x (1 + g) = 2 x (1 + 10%) = 2.20

Price per share of the common stock, P = 75

Recall Gordan Growth Model,

P = D1 / (Ke - g) where Ke = required return by shareholders = cost of equity

Hence, Ke = D1 / P + g = 2.20 / 75 + 10% = 12.93%

--------------------------

Compute the cost of preferred stock

Price per share of preferred stock, Ps = 80

Dividend, DS = 5

Recall Gordan Model with zero growth,

PS = DS / Kp where Kp = expected return on preferred stock = Cost of preferred stock

Hence, Kp = DS / PS = 5 / 80 = 6.25%

--------------------------------

Compute the Weighted Average Cost of Capital (WACC)

Target capital structure is:

  • Proportion of Common equity, We = 50%
  • Proportion of Preferred stock, Ws = 20% and
  • Proportion of Debt, Wd = 30%

hence, WACC = Wd x Kd + Ws x Ks + We x Ke ; Please note that in our case Kd represents post tax cost of debt and hence the factor of (1-T) is not required here.

WACC = 30% x 6.825% + 20% x 6.250% + 50% x 12.93% = 9.76%

------------------------

Using a WACC of 15%, apply four capital budgeting techniques to evaluate the project, assuming the Free Cash Flows are as follows: Years Free Cash Flows 0 $ -252,000.00 1 $118,625.00 2 $127,125.00 3 $181,000.00 The four techniques are NPV, IRR, MIRR, and discounted Payback. Assume the reinvestment rate to be 8% for the MIRR.

Please see the table below for NPV Calculation. The linkage column explains how each row has been calculated.

Year, N

Linkage

0

1

2

3

FCF

FCFF

(252,000)

118,625

127,125

181,000

WACC

R

15%

PV Factor

(1+R)(-N)

1.0000

0.8696

0.7561

0.6575

PV of FCF (Discounted cash flow)

FCF x PV Factor

(252,000)

103,152

96,125

119,010

NPV

Sum of all PVs

66,287

Since NPV is positive, the company should accept the project.

-----------------------------------

For the purpose of IRR, we will have to use the excel function of IRR.

IRR = IRR (C0, C1, C2,....Cn) = IRR (-252000, 118625,127125,181000) = 29.17%

IRR = 29.17% > WACC = 15%, the company should accept the project

---------------------

MIRR = (Future values of positive cash flows at reinvestment rate / PV of negative cash flows)(1/N) - 1 where N is the number of periods

Since there is only one negative cash flow in the project and that is the initial investment, the formula for MIRR gets simplified as, MIRR = (Future values of positive cash flows at reinvestment rate / Initial investment)(1/N) - 1

For Future values of positive cash flows at reinvestment rate, please see the table below:

Year, N

Linkage

1

2

3

Positive Cash flows

CF

118,625

127,125

181,000

Reinvestment rate

R

8%

Reinvestment period

T = 3 - N

2

1

0

FV factor

(1+R)^P

1.1664

1.08

1

FV of CF

138,364

137,295

181,000

Total FV of Positive cash flows

Sum of all FV of CF

456,659

MIRR = (456,659 / 252,000)(1/3) - 1 = 21.92%

Since MIRR = 21.92% > WACC = 15%, the company should accept the project.

---------------------------

Discounted Payback period

We have already calculated the discounted cash flow in the table above. Please see the table below now:

Year, N

Linkage

0

1

2

3

Discounted Cash flow

(252,000)

103,152

96,125

119,010

Cumulative discounted cash flow

Accumulate the discounted cash flow till that period

(252,000)

(148,848)

(52,723)

66,287

We can see that cumulative discounted cash flow assumes a zero value somewhere between year 2 and 3. We will have to interpolate between year 2 and 3.

Discounted payback period = 2 + 52,723 / 119,010 = 2.44 years

Criterion: The business will only accept projects with a payback period of two and half years or less. Since discounted payback period of 2.44 years is less than threshold of 2.50 years, the company should accept the project.

----------------------